CAT Chat Minutes

July 1, 2002

 

Information and Follow-up Items

1.   Lost mail causes problems. How can personnel be sure that they are not losing important e-mail?

The computer support group is intercepting over 1000 spam messages and 9 virus attacks a day. They review these each day to ensure that no legitimate messages are being intercepted. The spam assassin has been activated on @aps.anl.gov addresses only. The computer group filters all addresses for incoming viruses.

2.There is a proposal from ASD to use APS process water (de-ionized water) to cool beamline components. Gene Swetin will make a presentation at the July 18 TWG on this topic. What is the track history of clogging of APS process water-cooled components?

John Noonan, (ASD) answered that there had only been one such event very early in APS operations.

3.Is it necessary to have a Type C surveys for white and mono-beam mini-hutches?

This policy will be looked into at a BCRRT meeting next week. There is presently a standing order that requires a survey if white beam mini-hutches are secured after an access. For monobeam mini-hutches, the Floor Coordinator can use his/her discretion to permit beam.

User Operations

Operations Stats: Glen Decker reviewed the past two-week's operational events. Nine faults occurred since the last CAT Chat. Beam availability for week 5 is 94.31% and the mean time between faults is 39 hours.

The overall five-week period scheduled beam time is 675 hours. Nineteen faults have occurred so far leading to 19 hours of downtime. A twenty-four hour machine studies / intervention period will take place on Thursday, July 4 at 0800 hours. Operations will resume on July 5 at 0800 hours.

Multi-bunch Instability: Several of the faults have been connected to what appears to be a multi-bunch instability caused by a higher order mode in one of the 16 rf cavities. This will be further studied during upcoming studies.

Vacuum Event: Glen Decker mentioned that there was an accidental venting on the 12-ID beamline, which caused an EPS trip and beam dump. There was a hr delay before beam could be restored. Vacuum procedures exist and must be followed. Fortunately, there have only been two such trips in APS history.

Lattice Correction: The Accelerator Operations group is developing a push button operation to improve the lattice correction procedure after shutdowns. The lattice correction symmetrizes the lattice resulting in an improved lifetime and storage ring dynamic aperture. P.K. Job has developed a neutron detector for Sector 3-ID. This detector was installed and hooked up to the EPICS system for data logging. There has been a 3-fold decrease in neutron counts at 3-ID, which is consistent with the increased beam lifetime.

Higher Current Test: Glen displayed two viewgraphs of how the higher current test ran. At 10 am the current was ramped to 115 mA and from noon to 4 pm it was at approximately 125 mA Approximately 16ID lines had their gaps closed during this test. Roger Klaffky asked the group if any trouble occurred and if they would like to see this occur again in January 2003. No serious problems were noted, but more time would be required to really test beamline performance. There was agreement that another test should take place at approximately 110 mA and with no more frequent than two-minute top up intervals. Training will be provided to the operators on injection before the next high current test takes place. If a 110mA special operations one week period is added in January, there will have to be sufficient studies to ensure that the present top-off beam current stability is maintained so that beamline optics will be just as thermally stable. There was some preference for using the first week of operations in January for this study, but the CR will have to be ready to start up with this more demanding operational condition.

 

General Information

The Long Range Operations Schedule will be out towards the end of July adding the months of January through April.